OpenStudy (anonymous):

1) Solve the system of equations by graphing. Then classify the system x+y=10 x-y= -6

OpenStudy (anonymous):

Solution at: x = 2 y = 8

OpenStudy (anonymous):

Classification: consistent (unique solution)

OpenStudy (anonymous):

Oh and its a 2x2 system.

OpenStudy (anonymous):

to graph it, put both equations in terms of y: y = 10 - x y = 6 + x

OpenStudy (anonymous):

I posted a graph already.

OpenStudy (anonymous):

You didn't explain anything though

OpenStudy (anonymous):

True.

OpenStudy (anonymous):

It's called open study not open answers

OpenStudy (anonymous):

Then let me explain a bit further, once you write each equation as a function of x and graph the two functions, the solution to the system (if it exists) is where the two lines intersect. Furthermore if the two functions express the same line there are an infinite number of solutions.